# Open-Set Gait Recognition from Sparse mmWave Radar Point Clouds

### v2

@@ Abstract
The adoption of Millimeter-Wave (mmWave) radar devices for human sensing, particularly gait recognition, has recently gathered significant attention due to their efficiency, resilience to environmental conditions, and privacy-preserving nature. In this work, we tackle the challenging problem of Open-set Gait Recognition (OSGR) from sparse mmWave radar point clouds. Unlike most existing research, which assumes a closed-set scenario, our work considers the more realistic open-set case, where unknown subjects might be present at inference time, and should be correctly recognized by the system. Point clouds are well-suited for edge computing applications with resource constraints, but are more significantly affected by noise and random fluctuations than other representations, like the more common micro-Doppler signature. This is the first work addressing open-set gait recognition with sparse point cloud data. To do so, we propose a novel neural network architecture that combines supervised classification with unsupervised reconstruction of the point clouds, creating a robust, rich, and highly regularized latent space of gait features. To detect unknown subjects at inference time, we introduce a probabilistic novelty detection algorithm that leverages the structured latent space and offers a tunable trade-off between inference speed and prediction accuracy. Along with this paper, we release mmGait10, an original human gait dataset featuring over five hours of measurements from ten subjects, under varied walking modalities. Extensive experimental results show that our solution attains F1-Score improvements by $24\%$ over state-of-the-art methods, on average, and across multiple openness levels.
@@ I Introduction
To the best of our knowledge, no previous work in the literature jointly tackled both challenges, thus being unsuited for realistic edge-based people monitoring applications.
@@
    We acquire and publish mmGait10, a new high quality, multi-scenario, human-gait mmWave radar dataset, consisting of approximately $5$ hours of measurements. We release both the raw data and the pre-processed sparse point clouds of ten subjects freely walking around within an indoor environment with different walking modalities. The large bandwidth and high angular resolution of the employed mmWave device allowed for denser point clouds as compared to those available from existing datasets, with point clouds containing an average of $150$ points per frame.
@@
    We evaluate PCAA in different experimental conditions and compare it with the latest available methods from the literature. We demonstrate that our method significantly outperforms existing solutions, achieving an average improvement of $24\%$ in the F1111-Score.
The manuscript is structured as follows. In Section II, we provide a review of the existing literature on gait classification with mmWave radars, underlining the novel aspects of our work. In Section IV, we thoroughly describe our method and the PCAA architecture. In Section V, we provide details about the mmGait10 dataset, from the experimental methodologies to statistics about the captured point clouds. In Section VI, we show our experimental results. Concluding remarks are reported in Section VII.
@@ II-B Open-set gait recognition from radar measurements
    Latent space regularization: In closed-set classification, feature extraction is purely driven by supervised training to minimize the classification loss. Therefore, the resulting latent space of feature representations strongly separates samples associated with different labels. However, by only relying on the classification label information, the resulting distribution of feature representations will not be sufficiently representative of the general spatio-temporal structure of the input, which may be critical to OSGR. To discriminate among seen and unseen subjects, some form of unsupervised learning is necessary to extract descriptive features that contain more information than the classification label.
@@
Existing works tackle the above challenges using subject-specific $\mu$D spectrograms as input data.
@@
[21] presents a Deep Discriminative Representation Network (DDRN) trained with the cosine margin loss [22]. The latter promotes an increment of the inter-class distances and a reduction of the intra-class variance. The unknown detection is then performed by exploiting Extreme Value Theory (EVT), fitting the training feature vectors with a probability distribution to estimate the class inclusion probability. This approach allows finding a bound to the support region of the known classes in the embedding space, which is then used to recognize unknown subjects.
In [12], Open-set Recognition based on Contrastive constraint and Ensemble-based out-of-distribution Detection (OR-CED), a dual branch neural network based on the ladder Variational Auto-Encoders (VAE) architecture is proposed. OR-CED performs both reconstruction and classification of radar $\mu$D signatures. To counteract the negative effect of samples that are hard to discriminate, the authors introduce a contrastive loss term to force latent representations of each class to be close in the latent space. During inference, the distributions of the reconstruction errors and latent representations of training samples are modeled as univariate and multivariate Gaussians respectively. Then, two out-of-distribution detectors are jointly used to separate known and unknown subjects.
@@
The extraction of features from dynamic, sparse radar point clouds requires the adoption of specialized neural architectures capable of extracting features from individual multidimensional point clouds and their evolution in time. While neural architectures for $\mu$D signatures are largely taken from the vast image processing literature, advanced neural networks for radar point cloud sequences are less investigated.
@@
III Background on radar point clouds
@@
IV Proposed Method
In this work, we tackle the problem of OSGR from sparse radar point clouds. Our system consists of two main phases: deep feature extraction and unknown subject detection. In the first phase, the system extracts compact features of human gait from the input radar point cloud sequences. To this end, we propose PCAA, a deep neural network that reaps the benefits of supervised and unsupervised learning strategies. In Section IV-A we describe the pre-processing steps that are applied to the raw radar point clouds before they are fed to the network. Then, the network architecture and training strategy are thoroughly described in Section IV-B and Section IV-C respectively.
In the second phase, we leverage the regularized structure of the trained latent space to perform the novel subject detection task. To this end, we propose a simple detection algorithm, capable of determining whether an input gait feature vector belongs to a known or unknown subject, at inference time. The proposed novel subject detection algorithm is described in Section IV-D.
IV-A Point cloud pre-processing
@@
To tackle the OSGR task, we consider a dataset divided into a set of known subjects, to use for training, and a set of unknown subjects, never seen by the model during training. These are used to assess the ability of the network to recognize novel subjects. We define $S_{K}\subset S$ the set of known subject IDs and $S_{U}\subset S$ the set of the unknown ones, so that $\{S_{K},S_{U}\}$ is a partition of $S$. The cardinalities of $S_{K}$ and $S_{U}$ depend on the specific scenario. Following the OSGR literature, we measure how many unknown subjects are present with respect to the known ones using the openness metric, whose definition and impact on the performance of PCAA are presented in Section VI-A.
@@ IV-B Model architecture
Given the input tuple $(\mathbf{X}_{1:N_{f}},\mathbf{e}_{i})$, belonging to the known dataset partition $\mathcal{K}$, the key idea is to constrain the encoder to generate a representation $\mathbf{z}$ according to a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu}_{i},\mathbf{I}_{K})$, where $\mathbf{I}_{K}$ is the $K\times K$ identity matrix. In simpler terms, our primary aim is to shape the probability distribution induced by the encoder, with conditioning arising from label information, leading to samples of different classes being distributed in separate clusters in the latent space. This choice grants both good separability of the classes, to reach high classification accuracy in the closed set, and soft transitions between clusters, learning smooth semantic representations. The adversarial branch is only employed during training and is discarded during inference.
@@ IV-C Model training
IV-D Unkown subject detection
@@
Specifically, we exploit the fact that, after training, feature vectors extracted from radar traces of subject $i$ follow a normal distribution $\mathcal{N}(\boldsymbol{\mu}_{i},\mathbf{I}_{K})$, where the mean vectors $\{\boldsymbol{\mu}_{i}\}_{i=1}^{M}$ were deterministically chosen before training according to the procedure described in Section IV-B. We can therefore define the known distribution, as a mixture of Normal Gaussian distributions with equal weights, which is written as
@@
V mmGait10 dataset Structure
To evaluate our approach, we collect mmGait10, a radar point cloud dataset for human gait analysis. The dataset is available at https://zenodo.org/records/14974386. This comprises approximately 5 hours of radar point cloud measurements collected from $10$ subjects. The participants have a diverse range of physical characteristics, including sex, height, and weight, to capture a realistic variety of walking patterns. Informed consent was obtained from all the involved subjects before data collection. More details regarding the amount of recorded data for each subject are reported in Tab. 3. Each subject is instructed to walk freely along random trajectories in a $7.81$ m $\times$ $7.26$ m room. To further enhance diversity and realism, each subject’s gait is recorded under three different conditions: i) walking freely, ii) walking freely while holding a smartphone, and iii) walking freely with their hands in their pockets. These different walking modalities, where the motion of different body parts is present or absent (e.g., when holding a smartphone, the arm’s natural swinging motion is restricted), enable learning more general and robust gait features. Each walking modality is carried out for approximately $10$ minutes, for a total recording time of approximately $30$ minutes per subject. All the measurements are collected with a commercial Texas Instruments MMWCAS-RF-EVM FMCW Radar, operating in the $77$-$81$ GHz frequency band at a frame rate of $10$ Hz. We report the radar parameters and the resolution of our measurements in Tab. 2.
@@ VI-A Evaluation metrics
VI-B Comparison with existing methods
In this section, we compare PCAA to existing solutions in terms of predictive performance. As previously mentioned, there are currently no models specifically designed for solving the OSGR task using radar point cloud sequences. To provide a comparative evaluation with existing approaches, OR-CED [12] was selected as the benchmark, adapting it to our problem, as it is the most recent model in the literature addressing a similar task. OR-CED implements a probabilistic ladder architecture [32] [38] with a symmetric encoder-decoder structure, which was designed for the processing of micro-Doppler spectrograms and is not directly applicable to point cloud sequences. Therefore, for our comparisons, we implemented a variation of OR-CED which employs the same architectures of PCAA for its encoder and decoder modules. Hence, the comparison between the two models is focused on high-level architectural features rather than on the implementations of the individual modules. More specifically:
@@ VI-B Comparison with existing methods
    PCAA leverages the same approach of AAE, thus employing a dedicated adversarial module (the discriminator) to enforce a regularized distribution of the latent space. Conversely OR-CED uses a VAE architecture to achieve the same purpose, thus adopting an additional Kullback-Leibler divergence loss term.
@@
    OR-CED adopts an additional triplet loss term applied to vectors in latent space, which we do not use. This is because the desired clustering effect around the class centroids is intrinsically obtained by PCAA during the regularization phase.
@@
    While OR-CED learns the parameters of the class distributions in the latent space, with PCAA we fix the parameters and allow the network to focus on how the latent vectors should be arranged in the latent space, given the prior distribution.
We evaluate the performance of all models under different levels of openness to test their robustness to more challenging scenarios. We also report the performance of PCAA with different values of $k$ for the novel subject detection phase. Fig. 3 shows the performance of PCAA against OR-CED. We average our results over $n_{\text{test}}=5$ different trials, each employing a different dataset partitioning $(\mathcal{K},\mathcal{U})$. We report the mean F1111-score value along with a measure of the dispersion of the results, obtained as $\sigma/\sqrt{n_{\text{test}}}$, where $\sigma$ is the standard deviation of the F1111-scores over the $n_{\rm test}$ trials. We highlight that the randomization of the partitions is key to ensuring a fair evaluation. Indeed, the difficulty of the unknown detection is heavily influenced by which subjects are present in the unknown set, due to the individual variability of the gait features.
Figure 3: Macro-F1 scores of OR-CED and PCAA for different levels of openness. We also vary the number of aggregation steps $k$.
@@
It is important to take into account that OR-CED performs novel subject detection using a single input sample. Due to the high variability of human gait patterns, a single sample is sometimes not enough to differentiate among subjects. Conversely, our method is capable of waiting for $k\geq 1$ input samples before returning a prediction, observing a longer point could sequence to improve its confidence. To provide a fair comparison, we also report results for PCAA for $k=1$. It is clear that the complexity of the problem significantly increases with the openness for all the considered approaches. his causes a general decreasing trend in the F1111 score. However, for each openness level, our approach still manages to outperform the state of the art by a large amount. The large gap in F1111 score is due to the superior performance of our unknown subject detection method. Indeed, as shown in the confusion matrices in Fig. 6, OR-CED performs well in the closed set classification task but cannot discriminate between unknown and known targets. On the other hand, PCAA can distinguish the majority of unknown subjects, even using a single input sample (i.e. $k=1$).
Our results demonstrate that adapting existing approaches for OSGR to support sequential sparse radar point clouds is insufficient to achieve the same kind of performance. This further underscores the complexity of solving the OSGR task from radar point cloud sequences.
(a) OR-CED
(b) PCAA with $k=1$
(c) PCAA with $k=2$
(d) PCAA with $k=4$
Figure 6: Example of confusion matrices in a scenario with $21.55\%$ openness (9 subjects of which 4 known and 5 unknown). OR-CED is not able to detect novel samples, while PCAA can distinguish the majority of unknown subjects. The performances significantly improve as $k$ increases.
VI-C Ablation studies

### v4

@@ Abstract
The adoption of Millimeter-Wave (mmWave) radar devices for human sensing, particularly gait recognition, has recently gathered significant attention due to their efficiency, resilience to environmental conditions, and privacy-preserving nature. In this work, we tackle the challenging problem of Open-set Gait Recognition (OSGR) from sparse mmWave radar point clouds. Unlike most existing research, which assumes a closed-set scenario, our work considers the more realistic open-set case, where unknown subjects might be present at inference time, and should be correctly recognized by the system. Point clouds are well-suited for edge computing applications with resource constraints, but are more significantly affected by noise and random fluctuations than other representations, like the more common micro-Doppler signature. This is the first work addressing open-set gait recognition with sparse point cloud data. To do so, we propose a novel neural network architecture that combines supervised classification with unsupervised reconstruction of the point clouds, creating a robust, rich, and highly regularized latent space of gait features. To detect unknown subjects at inference time, we introduce a probabilistic novelty detection algorithm that leverages the structured latent space and offers a tunable trade-off between inference speed and prediction accuracy. Along with this paper, we release mmGait10, an original human gait dataset featuring over five hours of measurements from ten subjects, under varied walking modalities. Extensive experimental results show that our solution attains $24\%$ average F1-Score improvement over state-of-the-art methods adapted for point clouds, across multiple openness levels.
@@ I Introduction
To the best of our knowledge, no previous work in the literature has jointly tackled both challenges, thus being unsuited for realistic edge-based people monitoring applications.
@@
    We acquire and publish mmGait10, a new high quality, multi-scenario, human-gait mmWave radar dataset, consisting of approximately $5$ hours of measurements. We release both $\mu$D spectrograms and pre-processed sparse point clouds of ten subjects freely walking around within an indoor environment with different walking modalities. The large bandwidth and high angular resolution of the employed mmWave device allowed for denser point clouds as compared to those available from existing datasets, with point clouds containing an average of $150$ points per frame.
@@
    We evaluate PCAA in different experimental conditions and compare it with the latest available methods from the literature. We demonstrate that our method significantly outperforms existing solutions adapted for points clouds, achieving $24\%$ average F1111-Score improvement.
The manuscript is structured as follows. In Section II, we provide a review of the existing literature on gait classification with mmWave radars, underlining the novel aspects of our work. In Section III, we thoroughly describe our method and the PCAA architecture. In Section IV, we provide details about the mmGait10 dataset, from the experimental methodologies to statistics about the captured point clouds. In Section V, we show our experimental results. Concluding remarks are reported in Section VI.
@@ II-B Open-set gait recognition from radar measurements
    Latent space regularization: In closed-set classification, feature extraction is purely driven by supervised training to minimize the classification loss. Therefore, the resulting latent space of feature representations strongly separates samples associated with different labels. However, by only relying on the classification label information, the resulting distribution of feature representations will not be sufficiently representative of the general spatio-temporal structure of the input, which may be critical to OSGR. To discriminate among seen and unseen subjects, some form of strong regularization on the latent representations is needed, to achieve a structure which would not be possible to infer from the classification labels alone.
@@
Existing works tackle both challenges with a variety of approaches, ranging from generative models, contrastive learning strategies, or specialized loss functions.
@@
[21] presents a Deep Discriminative Representation Network (DDRN) trained with the cosine margin loss [22]. The latter promotes an increase in the inter-class distances and a reduction of the intra-class variance. The unknown detection is then performed by exploiting Extreme Value Theory (EVT), fitting the training feature vectors with a probability distribution to estimate the class inclusion probability. This approach allows finding a bound to the support region of the known classes in the embedding space, which is then used to recognize unknown subjects.
In a later work, [23], the same authors train an OSGR network through a Large margin Gaussian Mixture (L-GM) loss, which shapes the resulting latent space of gait features as a Mixture of Gaussians distribution where each subject is modeled as a different Gaussian component. By exploiting this property of the feature space, known users are directly classified according to the class-posterior probability, while unknown ones are identified by setting a probability threshold.
In [12], Open-set Recognition based on Contrastive constraint and Ensemble-based out-of-distribution Detection (OR-CED), a dual branch neural network based on the ladder Variational Auto-Encoders (VAE) architecture is proposed. OR-CED performs both reconstruction and classification of radar $\mu$D signatures. To counteract the negative effect of samples that are hard to discriminate, the authors introduce a contrastive loss term to force latent representations of each class to be close in the latent space. During inference, the distributions of the reconstruction errors and latent representations of training samples are modeled as univariate and multivariate Gaussians, respectively. Then, two out-of-distribution detectors are jointly used to separate known and unknown subjects.
@@
Extracting features from dynamic, sparse radar point clouds requires the adoption of specialized neural architectures capable of extracting features from individual multidimensional point clouds and their evolution in time. While neural architectures for $\mu$D signatures are largely taken from the vast image processing literature, advanced neural networks for radar point cloud sequences are less investigated.
@@
II-C Background on radar point clouds
@@
III Proposed Method
In this work, we tackle the problem of OSGR from sparse radar point clouds. Our system consists of two main phases: deep feature extraction and unknown subject detection. In the first phase, the system extracts compact features of human gait from the input radar point cloud sequences. To this end, we propose PCAA, a deep neural network that reaps the benefits of supervised and unsupervised learning strategies. In Section III-A we describe the pre-processing steps that are applied to the raw radar point clouds before they are fed to the network. Then, the network architecture and training strategy are thoroughly described in Section III-B and Section III-C respectively.
In the second phase, we leverage the regularized structure of the trained latent space to perform the novel subject detection task. To this end, we propose a simple detection algorithm, capable of determining whether an input gait feature vector belongs to a known or unknown subject, at inference time. The proposed novel subject detection algorithm is described in Section III-D.
III-A Point cloud pre-processing
@@
To tackle the OSGR task, we consider a dataset divided into a set of known subjects, to use for training, and a set of unknown subjects, never seen by the model during training. These are used to assess the ability of the network to recognize novel subjects. We define $S_{K}\subset S$ the set of known subject IDs and $S_{U}\subset S$ the set of the unknown ones, so that $\{S_{K},S_{U}\}$ is a partition of $S$. The cardinalities of $S_{K}$ and $S_{U}$ depend on the specific scenario. Following the OSGR literature, we measure how many unknown subjects are present with respect to the known ones using the openness metric, whose definition and impact on the performance of PCAA are presented in Section V-A.
@@ III-B Model architecture
Given the input tuple $(\mathbf{X}_{1:N_{f}},\mathbf{e}_{i})$, belonging to the known dataset partition $\mathcal{K}$, the key idea is to constrain the encoder to generate a representation $\mathbf{z}$ according to a Gaussian distribution $\mathcal{N}(\boldsymbol{\mu}_{i},\mathbf{I}_{K})$, where $\mathbf{I}_{K}$ is the $K\times K$ identity matrix. In simpler terms, our primary aim is to shape the probability distribution induced by the encoder, with conditioning arising from label information, leading to samples of different classes being distributed in separate clusters in the latent space. This choice grants both good separability of the classes, to reach high classification accuracy in the closed set, and soft transitions between clusters, learning smooth semantic representations. Notably, the adversarial and unsupervised branches are only employed during training and are discarded at inference time, ensuring a low memory footprint suitable for edge devices.
@@ III-C Model training
III-D Unkown subject detection
@@
Specifically, we exploit the fact that, after training, feature vectors extracted from radar traces of subject $i$ follow a normal distribution $\mathcal{N}(\boldsymbol{\mu}_{i},\mathbf{I}_{K})$, where the mean vectors $\{\boldsymbol{\mu}_{i}\}_{i=1}^{M}$ were deterministically chosen before training according to the procedure described in Section III-B. We can therefore define the known distribution, as a mixture of Normal Gaussian distributions with equal weights, which is written as
@@
IV mmGait10 dataset Structure
To evaluate our approach, we collect mmGait10, a radar dataset for human gait analysis [38], that includes both $\mu$D spectrograms and point cloud sequences. This comprises approximately 5 hours of radar measurements collected from $10$ subjects. The participants have a diverse range of physical characteristics, including sex, height, and weight, to capture a realistic variety of walking patterns. Informed consent was obtained from all the involved subjects before data collection. More details regarding the amount of recorded data for each subject are reported in Tab. 3. Each subject is instructed to walk freely along random trajectories in a $7.81$ m $\times$ $7.26$ m room. To further enhance diversity and realism, each subject’s gait is recorded under three different conditions: i) walking freely, ii) walking freely while holding a smartphone, and iii) walking freely with their hands in their pockets. These different walking modalities, where the motion of different body parts is present or absent (e.g., when holding a smartphone, the arm’s natural swinging motion is restricted), enable learning more general and robust gait features. Each walking modality is carried out for approximately $10$ minutes, for a total recording time of approximately $30$ minutes per subject. All the measurements are collected with a commercial Texas Instruments MMWCAS-RF-EVM FMCW Radar, operating in the $77$-$81$ GHz frequency band at a frame rate of $10$ Hz. We report the radar parameters and the resolution of our measurements in Tab. 2.
@@ V-A Evaluation metrics
V-B Performance evaluation
In this section, we compare PCAA to existing solutions in terms of predictive performance.
Methodology and choice of the baselines: To the best of our knowledge, our approach is the first to tackle the OSGR task from spatio-temporal point clouds. The lack of existing approaches for point-cloud-based OSGR poses a challenge in defining a fair and direct comparison with other established baselines. As a first step, we identify the most recent state-of-the-art approaches for OSGR, namely OR-CED [12] and L-GM [23], to use as baselines for our experiments. As discussed in Section II-B, these methods are designed for $\mu$D spectrograms. Therefore, to test these approaches on the same data modality for which they were designed, we also obtain $\mu$D spectrograms from our radar measurements and use them as training and testing data. However, to distinguish the impact of the different data representation ($\mu$D vs. point clouds) from that of the different architecture, we also test the performance of the baselines under the different data modality of point clouds, which is the main focus of this work. Indeed, the processing of point clouds entails additional challenges, such as temporal correlation modeling and ensuring the invariance of the output to input points permutations. Therefore, we decided to introduce a variation of each baseline approach, adapted for point-cloud processing. We will denote these two variants of OR-CED and L-GM as PC-OR-CED and PC-L-GM, respectively. We kept the modifications to the minimum needed to extend the baselines to point cloud data, to provide the fairest possible comparison. By testing the baselines both on spectrogram images and on spatio-temporal point clouds, we enable a comprehensive comparison that accounts both for architectural and data modality differences.
Besides the minimal needed changes, these point-cloud variants preserve the high-level architectural components of OR-CED and L-GM without changes, such as loss functions, regularization strategies, and overall network structure. More in detail:
  • The original OR-CED implements a probabilistic ladder architecture [33] [40] with a symmetric encoder-decoder structure, which was designed for the processing of micro-Doppler spectrograms and is not directly applicable to point cloud sequences. Therefore, in PC-OR-CED, we replace that part of the network with the Encoder and Decoder modules of PCAA, which are suitable for point clouds processing.
  • On the other hand, for PC-L-GM, we adapt the original approach by first reproducing the same network structure as in [23], but replacing the ResNet-18 feature extractor backbone, that is only suitable for image data, with the Encoder block of PCAA, similarly to what we do for PC-OR-CED. We then train the adapted network using the same original loss function.
Notably, our only modifications to the baselines concern the feature extraction network that operates on point clouds. Hence, our comparisons highlight high-level architectural features rather than the implementation of the individual modules, due to the different nature of the input data types involved (point cloud sequences and $\mu$D spectrograms).
More specifically:
@@ V-B Performance evaluation
    PCAA leverages the same approach of AAE, thus employing a dedicated adversarial module (the discriminator) to enforce a regularized distribution of the latent space. OR-CED and L-GM also aim at regularizing the latent space, but using different strategies, namely the VAE architecture and the L-GM loss function, respectively.
@@
    OR-CED adopts an additional contrastive loss term applied to vectors in latent space, which we do not use. This is because the desired clustering effect around the class centroids is intrinsically obtained by PCAA during the regularization phase.
@@
    While OR-CED and L-GM learn the parameters of the class distributions in the latent space, in PCAA we fix the parameters and allow the network to focus on how the latent vectors should be arranged in the latent space, given the prior distribution.
Results discussion: We evaluate the performance of all models under different levels of openness to test their robustness to more challenging scenarios. We also report the performance of PCAA with different values of $k$ for the novel subject detection phase. Fig. 3 shows the performance of PCAA against the benchmarks. We average our results over $n_{\text{test}}=5$ different trials, each employing a different dataset partitioning $(\mathcal{K},\mathcal{U})$. We report the mean F1111-score value along with a measure of the dispersion of the results, obtained as $\sigma/\sqrt{n_{\text{test}}}$, where $\sigma$ is the standard deviation of the F1111-scores over the $n_{\rm test}$ trials. We highlight that the randomization of the partitions is key to ensuring a fair evaluation. Indeed, the difficulty of the unknown detection is heavily influenced by which subjects are present in the unknown set, due to the individual variability of the gait features.
Additionally, our approach exhibits a clear trend of improved performance as the value of $k$ increases. This behavior represents one of the key advantages of our method over existing ones, as it enables users to balance inference speed and accuracy in the OSGR task by adjusting $k$ accordingly. For instance, in scenarios where only a few unknowns are expected, the openness of the problem can be assumed to be low, allowing small values of $k$ to yield both fast and accurate results. Conversely, in more challenging settings, where the number of unknowns is large, higher values of $k$ can be selected to reduce misclassifications at the cost of increased computational time.
Figure 3: OSGR macro-averaged F1 scores of OR-CED, L-GM, PCAA, PC-OR-CED and PC-L-GM, for different levels of openness. We also vary the number of aggregation steps $k$.
@@
It is important to take into account that both benchmarks perform novel subject detection using a single input sample. Due to the high variability of human gait patterns, a single sample is sometimes not enough to differentiate among subjects. Conversely, our method is capable of waiting for $k\geq 1$ input samples before returning a prediction, observing a longer point cloud sequence to improve its confidence. To provide a fair comparison, we also report results for PCAA for $k=1$. It is clear that the complexity of the problem significantly increases with the openness for all the considered approaches. This causes a general decreasing trend in the F1111 score.
Fig. 3 also shows the performance of the baselines, OR-CED and L-GM, trained and tested both with $\mu$D spectrogram data, as well as PC-OR-CED and PC-L-GM, trained and tested on our sequential point cloud data. There is a clear performance drop between L-GM and PC-L-GM, which shows that models designed for $\mu$D spectrograms are not suited for point cloud data. OR-CED performs poorly in both data representations. This could be attributed to the differences in data distribution between our dataset and the dataset employed in [12]. Notably, PCAA outperforms all the considered baselines, even when trained and tested employing their native data representations. To better showcase the performance of each considered approach, we report some examples of confusion matrices in Fig. 6
Our results demonstrate the superiority of PCAA over existing approaches on the task of OSGR from spatio-temporal point clouds. Our results also confirm that adapting existing methods for OSGR to support point clouds processing is insufficient to achieve the same level of performance. This further underscores the complexity of solving the OSGR task from radar point cloud sequences.
Additionally, to demonstrate the computational and memory efficiency of our approach, we report the observed metrics regarding training and inference times, and memory usage in Tab. 4.
TABLE 4: Time and memory requirements for PCAA. Training times are reported for different values of $S_{K}$, the number of subjects present in the training set. Inference times and input sizes are reported assuming input batches of $k$ samples, for different values of $k$.
Metric [unit] Case Value
Training time [min] $S_{K}=2$ $28.83\pm 0.66$
$S_{K}=4$ $57.15\pm 0.68$
$S_{K}=6$ $86.41\pm 2.02$
$S_{K}=8$ $119.58\pm 3.06$
Inference time [ms] $k=1$ $2.76\pm 0.43$
$k=2$ $5.10\pm 0.73$
$k=4$ $9.63\pm 1.31$
$k=6$ $14.02\pm 1.82$
Input size [MB] $k=1$ $0.069$
$k=2$ $0.137$
$k=4$ $0.275$
$k=6$ $0.412$
Model size [MB] Encoder $9.305$
Decoder* $821.155$
Discriminator* $0.017$
    Not used during inference
The reported values show that our approach exhibits negligible inference times, i.e., in the order of milliseconds, if compared to the actual duration of the input point cloud sequences, which include a few seconds of human gait information. This holds even for higher values of $k$. Moreover, PCAA has a remarkably low memory footprint, both in terms of model parameters and input point cloud representations. These features make our approach suitable for edge computing applications involving resource-constrained devices.
(a) PC-OR-CED
(b) PC-L-GM
(c) OR-CED [12]
(d) L-GM [23]
(e) PCAA with $k=1$
(f) PCAA with $k=2$
(g) PCAA with $k=4$
(h) PCAA with $k=6$
Figure 6: Example of confusion matrices in a scenario with $21.55\%$ openness (9 subjects of which 4 known and 5 unknown). OR-CED, PC-OR-CED and PC-L-GM perform very well on the closed-set recognition task, but are not able to detect novel samples. When trained and tested on spectrogram images, L-GM performs significantly better, but still often mistakes known subjects for unknowns. Conversely, PCAA can distinguish the majority of unknown subjects, and its performance significantly improves as $k$ increases.
V-C Ablation studies